OpenStudy (anonymous):

Medal help...I'm confused... A deck of cards consists of 8 blue cards and 5 white cards. A simple random sample (random draws without replacement) of 6 cards is selected. What is the chance that one of the colors appears twice as many times as the other?

OpenStudy (anonymous):

List the ways in which you could have (1) twice as many blue cards as white cards (e.g. 4 blue 2 white) and (2) twice as many white cards as blue cards (e.g. 4 white 2 blue). Then calculate the probability of each happening. Remember they're done without replacement, so after drawing the first card, there are only 12 cards left etc. And the order doesn't matter. So for 4 blue 2 white it's basically ((8*7*6*5)*(4*3))/(13*12*11*10*9*8). Or you can just find the number of ways it can happen and divide it by the total no. of possible outcomes. (13!)

OpenStudy (anonymous):

ok i'll try that...thanks

OpenStudy (anonymous):

w8...this is what i did before...13/13*7/12*6/11*5/10*5/9*4/8 that's 4 blue and 2 whites 13/13*4/12*3/11*2/10*8/9*7/8...that's or 4 whites and 2 blues

OpenStudy (anonymous):

and i got the answers for both...do i have to add it or multiply

OpenStudy (anonymous):

coz i got .0442 on 1 combination and .01414 on the other

OpenStudy (zarkon):

you are putting an order on the blue and white cards by doing it the way you are

OpenStudy (anonymous):

what do you mean by that...i don't get it

OpenStudy (zarkon):

for the 4 blue and 2 white ... are you looking to get the first 4 blue and the second 2 white or do you just want 4 blue and 2 white?

OpenStudy (anonymous):

just 4 blue and 2 whites

OpenStudy (zarkon):

then you need to modify how you are computing the answer

OpenStudy (anonymous):

thank you

OpenStudy (zarkon):

I would use \({n\choose r}\) to help compute the answer

OpenStudy (zarkon):

or \(_nC_r\) if you are not familiar with the above notation

OpenStudy (zarkon):

so if you just wanted 4 blue and 2 white you would do this \[\frac{{8\choose 4}{5\choose 2}}{{13\choose 6}}\]

OpenStudy (anonymous):

I'm having problem with prob calculator ( I dont know why maybe because o the laguague barrier..i really dont know) that's why I'm doing most of the problem the long way

OpenStudy (zarkon):

you can do it the long way but you have to be mindful of accidentally putting an order on the section process...it is very easy to mess up.

OpenStudy (zarkon):

if you have \[\frac{(8\times7\times6\times5)\times(5\times4)}{13\times12\times11\times10\times9\times8}\] then you have put an order on the cards

OpenStudy (zarkon):

ie you just fond the prob of getting in order b,b,b,b,w,w

OpenStudy (zarkon):

there are \[\frac{6!}{4!2!}\] ways to permute the b's and w's

OpenStudy (zarkon):

so your answer (atlest for 4 blue and 2 white ) is \[\frac{(8\times7\times6\times5)\times(5\times4)}{13\times12\times11\times10\times9\times8}\times\frac{6!}{4!2!}\]

OpenStudy (zarkon):

using shorthand...that is \[\frac{{8\choose 4}{5\choose 2}}{{13\choose 6}}\]

OpenStudy (anonymous):

do i need to get the prob of both combination or just choose 1 combination

OpenStudy (zarkon):

you need to calculate P(4B and 2W)+P(2B and 4W) if that is what you are asking

OpenStudy (anonymous):

ok..got it..thanks a lot

OpenStudy (anonymous):

Zarkon...can you tell me if I did this right this time... I roll a die repeatedly. Find the chance that the first 4 rolls all show different faces, and the 5th roll shows a face that has appeared before. 6/6*5/6*4/6*3/6*4/6